OpenStudy (anonymous):

help please: find particular solution: cos^2(y^2)cos(x)dx-2sin(x)ydy=0 y(pi/2)=sqr((pi/2))

OpenStudy (jamesj):

To confirm: \[\cos^2 y^2 . \cos x. dx - 2 \sin x . y . dy = 0\]

OpenStudy (jamesj):

is that right?

OpenStudy (jamesj):

you there?

OpenStudy (jamesj):

Ok ... I see how to solve it. Let me know when you're back.

OpenStudy (anonymous):

Im here, is right

OpenStudy (jamesj):

Ok, the first thing you want to get rid of are the y^2 terms. So substitute u = y^2 and this will simplfy the equation somewhat. Write down what you get in terms of u and x.

OpenStudy (jamesj):

Yes we have the same thing. Now follow the first step I just gave you.

OpenStudy (anonymous):

and with 2sinxy?

OpenStudy (jamesj):

If u = y^2, then du = 2y dy.

OpenStudy (jamesj):

So cos^2(y^2)cos(x)dx-2sin(x)ydy=0 becomes cos^2 u . cos x . dx - sin x du = 0 Now you should have no problem in theory solving that equation because it is separable.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nicely done jj. Medal for you.

OpenStudy (anonymous):

cos^2u/du=sinx/cosx ·dx

OpenStudy (jamesj):

doesn't look right.

OpenStudy (anonymous):

How I do separate rightly?

OpenStudy (jamesj):

cos^2 u . cos x . dx - sin x du = 0 => cos^2 u . cos x . dx = sin x du => cos x/sin x . dx = sec^2 u du

OpenStudy (anonymous):

ok, them integrate?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

log(sin(X)) =tan(u) du ?

OpenStudy (anonymous):

log(sin(X)) =tan(u) du ?

OpenStudy (jamesj):

ln(sin x) = tan u + C, yes

OpenStudy (jamesj):

Well, first I'd solve for y(x) and then find the particular solution.

OpenStudy (anonymous):

and now to find the particular solution ?

OpenStudy (anonymous):

undo the u and replace the value, right?

OpenStudy (jamesj):

solve for y, yes

OpenStudy (anonymous):

thank u very much 4 your help!